OpenStudy (anonymous):

Find the derivative of f(x) = -12x^2 + 9x at x = 6.

OpenStudy (anonymous):

-112.5 -135 -90 -108

OpenStudy (epoweritheta):

f'=-24x+9

OpenStudy (anonymous):

f(x) = -12x^2 + 9x f(6) = -12(6)^2 + 9(6) = -12 (36) + 54 = -432 + 54 = -378

OpenStudy (anonymous):

^That's what I got but it's not an answer choice

OpenStudy (anonymous):

but thats the correct one

OpenStudy (anonymous):

I know

OpenStudy (johnweldon1993):

No that is the process of simply putting in 6 for 'x' you need first to find the derivative of the function... \[\large f(x) = -12x^2 + 9x\] \[\large f'(x) = -24x + 9\] and you THEN substitute the value of 6 in for 'x'

OpenStudy (anonymous):

oh ya then u'll get -135

OpenStudy (anonymous):

thanks guys!!

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (anonymous):

welcome